# An empirical study on code review activity prediction in practice

### v1

@@ 4.4.2. Large-Language Model encoding
To work with LLM models, the Hugging Face community provides many pretrained models available to use as-is (huggingface, ). We selected Bloom from Big Science, an LLM trained on many programming languages (e.g., Java, C++, Python), who’s direct use is text generation and exploring charateristics of language (bloom, ). The advantage of Bloom is that it provides a condensed version of the model that still provides good results for NLP tasks. We do use big science/bloom-560m in the context of this paper, which is about 4 GB compared to the normal version of Bloom, which is about 375 GB. Also, it counts more than 1.4M downloads in the month of August 2023, making it the most popular LLM pre-trained on programming languages at the time of our empirical study. To obtain the encoding, we used SentenceTransformer, a Hugging Face Python library (sentencetrans, ). For the encoding step, we do not clean the code, as Bloom is trained on raw code.
@@ 4.4.3. Review process features
software tasks such as risk prediction (nayrolles2018clever, ; shihab2012industrial, ) and bug prediction (giger2012method, ; kamei2012large, ) have been shown to use software process features for their training and getting respectable performance. Based on their example and inspired by their features, we defined review process features following two main categories: size-related features and historical features.
@@ 4.5. Classification
Similar to Hong et al. (hong2022should, ), we considered Random Forest (RF), Naive Bayes (NB), Nearest Neighbors, Logistic Regression, Decision Trees, and off-the-shelf LLMs (bloom560m) to classify the changed files as ”revised”, ”commented” and/or ”hot-spots” as three separated tasks. From preliminary analysis on those classifiers, we narrowed down the list to Random Forest and Naive Bayes, as they performed better in terms of AUC and F1-score for our task, regardless of the used embeddings. This confirmed Hong et al.’s observation (hong2022should, ) that Random Forest gives the best AUC for their task, followed in most cases by Naive Bayes. During preliminary analysis, we used the off-the-shelf Bloom LLM classifier, yet this did not perform as well as the considered ML models. As all the datasets are imbalanced considering the changed file granularity, we consider the SMOTE data sampling approach during training (wang2006classification, ; chawla2002smote, ). Models are trained on a per-project basis.
@@ 4.6. Evaluation
We empirically evaluate the models for the three tasks using three open-source and two closed-source datasets. To do the evaluation, we report AUC, as Hong et al. (hong2022should, ). We also report metrics used for imbalanced labeled datasets: F1-score (F1 = $\frac{2}{pre^{-1}+rec^{-1}}$) and Geometric Mean (GM = $\sqrt{pre\cdot rec}$), which are computed based on the precision (pre = $\frac{TP}{TP+FP}$) and recall (rec = $\frac{TP}{TP+FN}$). Those metrics need a threshold to define when the prediction is considered as true or false. We choose a threshold that maximizes the F1-score on the validation sets, as it captures best imbalance of the datasets (chase2014thresholding, ).
@@ 4.7. Statistical tests
Since we have a limited number of runs, we decided to group the runs into sub-categories of datasets: open-source datasets (i.e., Base, Ironic and Nova), closed-source datasets (i.e., Closed 1 and 2) and all datasets. When comparing two setups for a given label, we take all the runs of that setup for all the datasets in the subcategory. We apply a Wilcoxon paired test (cuzick1985wilcoxon, ) with $\alpha$-value = .01. We calculate the Cliff’s Delta effect to evaluate the impact of the significant differences (macbeth2011cliff, ).
@@ 5.1. Motivation
In their paper, Hong et al. explore the use of the Bag-of-Words (BoW) to embed code changes (hong2022should, ). Contrary to previous work using full-code changes to perform prediction (wattanakriengkrai2020predicting, ), they only consider added lines of a code change. This choice has the advantage of not having removed or unchanged lines contaminating the text embedding, as those lines could appear as noise in the training. However, the prediction will thus have less context since only the added lines are considered.
@@ 5.2. Approach
In RQ1, we explore how text embeddings impact the review activity predictions at file-level. Following the findings of Hong et al. (hong2022should, ), we will explore both (1) keeping only the added lines of a code change and (2) consider embedding removed lines in addition to the add lines. Among the different variations of each text embedding, we will identify the best embedding for a given dataset and label as ”Best text” embedding.
@@ 5.3. Results
For BoW, we found that 95-99% of the feature importance is focused on 1-5% of top features, compared to 50% of feature importance focused on 20-40% of top features for the LLM embedding. We apply a feature importance analysis using Python’s feature_importances_, which measures the percentage of feature importance of each individual feature. Comparison of the ”BoW +” and ”LLM +” embeddings shows that due to the very large size of the embedding outputs of BoW, the classifier will only focus on a low number of features, with the majority of stored features not playing an important role in the prediction. More details on feature importance can be found in the replication package (replic, ).
@@ 6.1. Motivation
The state-of-the-art approach only considers two basic software process features, i.e., number of lines add and removed, which were able to improve the performance of commented and revised prediction tasks (hong2022should, ). We thus want to evaluate how review process features in general, such as size-related features, but also features based on the history of the file, the author and the reviewers, perform compared to text embeddings, inspired by other software processes that have used similar features in the past (shihab2012industrial, ; kamei2012large, ).
@@ 6.2. Approach
  • ”Add/Rem”: this set of features refers to the features used by Hong et al. in their approach (hong2022should, ), i.e., ”c_add” and ”c_rem” in Table 5.
@@ 6.3. Results
Through these results, we observe that in the case of open-source datasets, the review process feature-based approach does not consistently perform as well as the ”Best text” embedding, though the F1-score reported by the best feature embeddings per dataset and label improved performance in 4/9 cases (up to +12%). Regarding closed-source datasets, the ”All feat” embedding performs as well as the ”Best text” embedding, while reducing the computational effort of model training thanks to the reduced dimensionality. Looking at feature importance per feature category of Table 5, we found that all datasets benefit most from the ”Add/Rem” and ”Count” categories, whose features each represent 2 to 13% of the overall feature importance. This indicates that those features regarding the size of changed files have the most value, including also the contextual features regarding the full patch size. More details on feature importance can be found in the replication package (replic, ).
@@ 7.1. Motivation
Hong et al. propose in their paper a review activity prediction approach using a BoW embedding in addition to features measuring the number of lines added and removed to a file (hong2022commentfinder, ). In RQ1 and RQ2, we explored how text embeddings and review process feature embeddings perform separately on the three tasks and what they capture about the code submission. In this question, we combine both types of embeddings to find the best embedding, compared to the state-of-the-art.
@@ 7.2. Approach
We combine each text embedding with each review process feature embedding presented in the previous RQs. The results are shown in Table 8. Each project is presented with three comparisons: the baseline (state-of-the-art (hong2022should, ), i.e., BoW embedding with ”Add/Rem” features), ”Best text” (from RQ1), and ”Best feat” (from RQ2). We report two main combinations: first, ”Heavy combo” which refers to our most ”heaviest” embedding compared to the baseline, i.e., the one considering the ”LLM +/-” text embedding and ”All feat” feature-based embedding. Then, we also report ”Best combo”, which reports the combination with the best F1-score per dataset and label. The lines directly following ”Best combo” report the specific combination chosen (i.e., the name of the text and review process feature embeddings used).
@@ 8.2. Computation time
One of our motivations for evaluating embeddings is that past work reports that simpler approaches sometimes work better in the case of software engineering prediction tasks (hong2022commentfinder, ). The use of feature-based approaches, and thus low dimensionality representations, has the advantage of low storage complexity and lower computational effort for training. Text embedding, on the other hand, depending on the approach, can be quite expensive to compute and store, and also increases the training effort. For the example of BoW, the size of the vocabulary explodes with the size of the corpus increasing. LLM encoding approaches usually embed the text into lower dimensions compared to BoW (e.g., around 2k for Bloom).
@@ 9. Threats to Validity
We used a zero-shot LLM-encoder as to reduce the computational effort. Fine-tuning the LLMs, though being computationally expensive, could improve the embedding and thus the results. However, past work has studied LLM fine-tuning for ”commented” prediction (li2022automating, ) and reported similar results to other Bag-of-Words embeddings (hong2022should, ). Larger LLMs might lead to better embeddings, even though the smaller bloom-560m has been reported as more performant than bigger counterparts in some other use-cases (sharma2023argumentative, ), hence it is not uncommon in research to leverage smaller models. As a compromise, we considered bloom-560m as a lower bound of the LLM embedding approach.
We worked on a per-project basis for embedding, training and evaluation, which is the default approach in this domain (hong2022commentfinder, ; li2022automating, ; tufano2022using, ). An early bootstrap option, when data is not yet available, might be to initially train on similar datasets (e.g., same coding languages).
@@ 10. Conclusion
In this paper, we explored different approaches for embedding code changes to predict review activities. This was motivated by the results of a user study involving 29 professional developers showing how ordering files in patches by prioritizing files that need review activity improves the review process. Indeed, we observe that such ordering approach lead to more review comments (+23%), that targeted parts needing revisions or comments (+13% precision and +8% recall), thus the need for an efficient review automation prediction approach. While previous work used a Bag-of-Word approach on the added lines and additional features (i.e., #lines added and #lines removed) (hong2022should, ), we proposed variations to that embedding.

### v2

@@ 4.4.2. Large-Language Model encoding
To work with LLM models, the Hugging Face community provides many pretrained models available to use as-is (hug, [n. d.]). We selected Bloom from Big Science, an LLM trained on many programming languages (e.g., Java, C++, Python), who’s direct use is text generation and exploring charateristics of language (BigScience, 2022). The advantage of Bloom is that it provides a condensed version of the model that still provides good results for NLP tasks. We do use big science/bloom-560m in the context of this paper, which is about 4 GB compared to the normal version of Bloom, which is about 375 GB. Also, it counts more than 1.4M downloads in the month of August 2023, making it the most popular LLM pre-trained on programming languages at the time of our empirical study. To obtain the encoding, we used SentenceTransformer, a Hugging Face Python library (sen, [n. d.]). For the encoding step, we do not clean the code, as Bloom is trained on raw code.
@@ 4.4.3. Review process features
software tasks such as risk prediction (Nayrolles and Hamou-Lhadj, 2018; Shihab et al., 2012) and bug prediction (Giger et al., 2012; Kamei et al., 2012) have been shown to use software process features for their training and getting respectable performance. Based on their example and inspired by their features, we defined review process features following two main categories: size-related features and historical features.
@@ 4.5. Classification
Similar to Hong et al. (Hong et al., 2022a), we considered Random Forest (RF), Naive Bayes (NB), Nearest Neighbors, Logistic Regression, Decision Trees, and off-the-shelf LLMs (bloom560m) to classify the changed files as ”revised”, ”commented” and/or ”hot-spots” as three separated tasks. From preliminary analysis on those classifiers, we narrowed down the list to Random Forest and Naive Bayes, as they performed better in terms of AUC and F1-score for our task, regardless of the used embeddings. This confirmed Hong et al.’s observation (Hong et al., 2022a) that Random Forest gives the best AUC for their task, followed in most cases by Naive Bayes. During preliminary analysis, we used the off-the-shelf Bloom LLM classifier, yet this did not perform as well as the considered ML models. As all the datasets are imbalanced considering the changed file granularity, we consider the SMOTE data sampling approach during training (Wang et al., 2006; Chawla et al., 2002). Models are trained on a per-project basis.
@@ 4.6. Evaluation
We empirically evaluate the models for the three tasks using three open-source and two closed-source datasets. To do the evaluation, we report AUC, as Hong et al. (Hong et al., 2022a). We also report metrics used for imbalanced labeled datasets: F1-score (F1 = $\frac{2}{pre^{-1}+rec^{-1}}$) and Geometric Mean (GM = $\sqrt{pre\cdot rec}$), which are computed based on the precision (pre = $\frac{TP}{TP+FP}$) and recall (rec = $\frac{TP}{TP+FN}$). Those metrics need a threshold to define when the prediction is considered as true or false. We choose a threshold that maximizes the F1-score on the validation sets, as it captures best imbalance of the datasets (Chase Lipton et al., 2014).
@@ 4.7. Statistical tests
Since we have a limited number of runs, we decided to group the runs into sub-categories of datasets: open-source datasets (i.e., Base, Ironic and Nova), closed-source datasets (i.e., Closed 1 and 2) and all datasets. When comparing two setups for a given label, we take all the runs of that setup for all the datasets in the subcategory. We apply a Wilcoxon paired test (Cuzick, 1985) with $\alpha$-value = .01. We calculate the Cliff’s Delta effect to evaluate the impact of the significant differences (Macbeth et al., 2011).
@@ 5.1. Motivation
In their paper, Hong et al. explore the use of the Bag-of-Words (BoW) to embed code changes (Hong et al., 2022a). Contrary to previous work using full-code changes to perform prediction (Wattanakriengkrai et al., 2020), they only consider added lines of a code change. This choice has the advantage of not having removed or unchanged lines contaminating the text embedding, as those lines could appear as noise in the training. However, the prediction will thus have less context since only the added lines are considered.
@@ 5.2. Approach
In RQ1, we explore how text embeddings impact the review activity predictions at file-level. Following the findings of Hong et al. (Hong et al., 2022a), we will explore both (1) keeping only the added lines of a code change and (2) consider embedding removed lines in addition to the add lines. Among the different variations of each text embedding, we will identify the best embedding for a given dataset and label as ”Best text” embedding.
@@ 5.3. Results
For BoW, we found that 95-99% of the feature importance is focused on 1-5% of top features, compared to 50% of feature importance focused on 20-40% of top features for the LLM embedding. We apply a feature importance analysis using Python’s feature_importances_, which measures the percentage of feature importance of each individual feature. Comparison of the ”BoW +” and ”LLM +” embeddings shows that due to the very large size of the embedding outputs of BoW, the classifier will only focus on a low number of features, with the majority of stored features not playing an important role in the prediction. More details on feature importance can be found in the replication package (rep, [n. d.]).
@@ 6.1. Motivation
The state-of-the-art approach only considers two basic software process features, i.e., number of lines add and removed, which were able to improve the performance of commented and revised prediction tasks (Hong et al., 2022a). We thus want to evaluate how review process features in general, such as size-related features, but also features based on the history of the file, the author and the reviewers, perform compared to text embeddings, inspired by other software processes that have used similar features in the past (Shihab et al., 2012; Kamei et al., 2012).
@@ 6.2. Approach
  • ”Add/Rem”: this set of features refers to the features used by Hong et al. in their approach (Hong et al., 2022a), i.e., ”c_add” and ”c_rem” in Table 5.
@@ 6.3. Results
Through these results, we observe that in the case of open-source datasets, the review process feature-based approach does not consistently perform as well as the ”Best text” embedding, though the F1-score reported by the best feature embeddings per dataset and label improved performance in 4/9 cases (up to +12%). Regarding closed-source datasets, the ”All feat” embedding performs as well as the ”Best text” embedding, while reducing the computational effort of model training thanks to the reduced dimensionality. Looking at feature importance per feature category of Table 5, we found that all datasets benefit most from the ”Add/Rem” and ”Count” categories, whose features each represent 2 to 13% of the overall feature importance. This indicates that those features regarding the size of changed files have the most value, including also the contextual features regarding the full patch size. More details on feature importance can be found in the replication package (rep, [n. d.]).
@@ 7.1. Motivation
Hong et al. propose in their paper a review activity prediction approach using a BoW embedding in addition to features measuring the number of lines added and removed to a file (Hong et al., 2022b). In RQ1 and RQ2, we explored how text embeddings and review process feature embeddings perform separately on the three tasks and what they capture about the code submission. In this question, we combine both types of embeddings to find the best embedding, compared to the state-of-the-art.
@@ 7.2. Approach
We combine each text embedding with each review process feature embedding presented in the previous RQs. The results are shown in Table 8. Each project is presented with three comparisons: the baseline (state-of-the-art (Hong et al., 2022a), i.e., BoW embedding with ”Add/Rem” features), ”Best text” (from RQ1), and ”Best feat” (from RQ2). We report two main combinations: first, ”Heavy combo” which refers to our most ”heaviest” embedding compared to the baseline, i.e., the one considering the ”LLM +/-” text embedding and ”All feat” feature-based embedding. Then, we also report ”Best combo”, which reports the combination with the best F1-score per dataset and label. The lines directly following ”Best combo” report the specific combination chosen (i.e., the name of the text and review process feature embeddings used).
@@ 8.2. Computation time
One of our motivations for evaluating embeddings is that past work reports that simpler approaches sometimes work better in the case of software engineering prediction tasks (Hong et al., 2022b). The use of feature-based approaches, and thus low dimensionality representations, has the advantage of low storage complexity and lower computational effort for training. Text embedding, on the other hand, depending on the approach, can be quite expensive to compute and store, and also increases the training effort. For the example of BoW, the size of the vocabulary explodes with the size of the corpus increasing. LLM encoding approaches usually embed the text into lower dimensions compared to BoW (e.g., around 2k for Bloom).
@@ 9. Threats to Validity
We used a zero-shot LLM-encoder as to reduce the computational effort. Fine-tuning the LLMs, though being computationally expensive, could improve the embedding and thus the results. However, past work has studied LLM fine-tuning for ”commented” prediction (Li et al., 2022) and reported similar results to other Bag-of-Words embeddings (Hong et al., 2022a). Larger LLMs might lead to better embeddings, even though the smaller bloom-560m has been reported as more performant than bigger counterparts in some other use-cases (Sharma et al., 2023), hence it is not uncommon in research to leverage smaller models. As a compromise, we considered bloom-560m as a lower bound of the LLM embedding approach.
We worked on a per-project basis for embedding, training and evaluation, which is the default approach in this domain (Hong et al., 2022b; Li et al., 2022; Tufano et al., 2022). An early bootstrap option, when data is not yet available, might be to initially train on similar datasets (e.g., same coding languages).
@@ 10. Conclusion
In this paper, we explored different approaches for embedding code changes to predict review activities. This was motivated by the results of a user study involving 29 professional developers showing how ordering files in patches by prioritizing files that need review activity improves the review process. Indeed, we observe that such ordering approach lead to more review comments (+23%), that targeted parts needing revisions or comments (+13% precision and +8% recall), thus the need for an efficient review automation prediction approach. While previous work used a Bag-of-Word approach on the added lines and additional features (i.e., #lines added and #lines removed) (Hong et al., 2022a), we proposed variations to that embedding.